OpenStudy (anonymous):

Using the following equation, find the center and radius of the circle. x2 + 4x + y2 − 6y = −4

OpenStudy (anonymous):

@Jennjuniper can you help?

OpenStudy (anonymous):

@mipie13 can you help?

OpenStudy (anonymous):

@cwrw238 can you help?

OpenStudy (anonymous):

The attached plot may give you some ideas.

OpenStudy (anonymous):

i will take a look at it!

OpenStudy (anonymous):

i need to know how to find the answer with the equation

OpenStudy (anonymous):

Organise the x and y terms so that they are aligned together so that it is easier to understand. Also subtract 3 from both sides: x^2 - 4x + y^2 + 6y = -3 Complete the square for each set of terms: x^2 - 4x + 4 + y^2 + 6y + 9 = 10 Factor: (x - 2)^2 + (y + 3)^2 = 10 The centre of a circle is at (0,0) and the radius is the square root of the result, thus: Centre = (2, -3) Radius = √10

OpenStudy (anonymous):

Centre = (2, -3) Radius = √10

OpenStudy (anonymous):

1. (x^2+4x+4)-4+(y^2-6y+9)-9=-4 (x+2)^2+(y-3)^2=9 (x+2)^2+(y-3)^2=3^2 Circles follow the form: (x-a)+(y-b)=r^2 Therefore the center of the circle is at (-2,3) and the radius is 3.

OpenStudy (anonymous):

thank you both :D i see what i wasnt understanding at first but using your responces i can figure out these problems now

OpenStudy (anonymous):

try google and type your question in that's what I normally do